Marshall Company is issuing four-year bonds with a coupon rate of 4.5 percent and semiannual coupon payments. If the current market rate for similar bonds is 9.8 percent, what will be the bond price? Round to 2 decimal places.
Answer #1

Price of a bond is mathematically represented as: 1 1 (1+ i) P C i (1) where P is price of bond, with periodic coupon C, n p

M = $1000, n = 4 * 2 = 8 semi-annual periods, C = 4.5% * $1000/2 = $22.50, i = 9.8%/2 = 4.9% (semi-annually)

1 (1+0.049)s 1000 P 22.50 (0.049)8 0.049

P = $146.01 + $682.02

P = $828.03